# Local converse theorems and Langlands parameters

### v1

@@ 6. A local converse theorem for Langlands parameters
Let $G$ be a quasi-split group with acceptable Langlands dual group. For $i=1,2$, let $\phi_{i}:{\mathrm{WD}}_{F}\to{}^{L}G$ be a generic homomorphism.
@@
    If $G$ is semi-simple and simply connected, it is sufficient to take $r$ above amongst the finite number of fundamental representations of ${}^{L}G$, and to twist by irreducible representations $\phi$ of ${\mathrm{WD}}_{K}$ of dimension at most $d$, where $d:=\lfloor N/2\rfloor$ for $N$ the dimension of any fundamental representation of highest dimension.
@@
Conjecture 6.1 is alslo true for non split but quasi-split orthogonal groups. See Theorem 7.2 below.
@@ 7. A variant
Statement 7.1.
Let $G=\mathrm{G}(F)$ be a quasi-split group with acceptable Langlands dual group. Then there exists $K/F$ a finite extension and $r:{}^{L}\mathrm{G}(K)\to{\mathrm{GL}}(V)$ a faithful representation of dimension $N$, such that if $\phi_{i}:{\mathrm{WD}}_{F}\to{}^{L}G$ is a generic homomorphism for $i=1,2$, and
@@
Statement 7.1 is true for the following list of $G$, $K$ and $r$.
@@
    $G={\mathrm{GL}}_{n}(F)$, $K=F$ and $r:{\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}\to{\mathrm{GL}}_{n}({\mathbb{C}})$ is the identity on ${\mathrm{GL}}_{n}({\mathbb{C}})$ extended trivially on $\mathrm{W}_{F}$, or $G={\mathrm{U}}_{n}(E/F)$, $K=E$, and $r:{\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{E}\to{\mathrm{GL}}_{n}({\mathbb{C}})$ is the identity on ${\mathrm{GL}}_{n}({\mathbb{C}})$ extended trivially on $\mathrm{W}_{E}$.
@@
The statement is obviously true in the first subcase of Case (a). For unitary groups it follows from the fact that $\phi_{1}\sim\phi_{2}$ if and only if ${\phi_{1}}_{|{\mathrm{WD}}_{E}}\sim{\phi_{2}}_{|{\mathrm{WD}}_{E}}$ as is well-known (see for example [Pra, Proposition 7.10]). In Case (b), the normalizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is equal to ×${\mathbb{C}}^{\times}\mathrm{G}_{2}({\mathbb{C}})$. This follows for example from applying [Gri, Theorem 1.1], and observing that the centralizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is the center of ${\mathrm{GL}}_{7}({\mathbb{C}})$ according to Schur’s lemma. Then the theorem in Case (b) follows Lemma 2.1, Theorem 3.3 and [Gri, Theorem 1.1]. In cases (c), (d), (e) and (f), the normalizers are respectively ×${\mathbb{C}}^{\times}{\mathrm{SO}}_{2n+1}({\mathbb{C}})$, ×${\mathbb{C}}^{\times}{\mathrm{Sp}}_{2n}({\mathbb{C}})$ and ×${\mathbb{C}}^{\times}{\mathrm{O}}_{2n+2}({\mathbb{C}})$, as follows for example from [Gri, Section 2]. The result that we claim follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 2.3]. ∎
@@
To conclude this paper, we make the following observation for classical groups when $F$ has characteristic zero. Assume the validity of the Gross-Prasad and Rallis conjecture and that the generic local Langlands correspondence $i$ of [JL3] coincides with that obtained by endoscopic transfer in [Art] and [Mok]. This latter fact might be known, and we refer to [AHKO, Theorem B.2 and Remark B.3] in the case of symplectic and split orthogonal case. Then using [JL3, Theorem 6.20] and thanks to the uniqueness of generic representations inside L-packets with respect to a prescribed non degenerate character proved in [Ato], Theorem 7.2 coincides with the local converse theorems mentioned in the introduction, using Langlands-Shahidi gamma factors.

### v4

@@ 6. A local converse theorem for Langlands parameters
Let $G$ be a quasi-split group with acceptable Langlands dual group. For $i=1,2$, let $\phi_{i}:\mathrm{W}_{F}\to{}^{L}G$ be an admissible homomorphism with associated Weil-Deligne parameter $(\varphi_{i},N_{i})$.
@@
    If $G$ is $F$-split, semi-simple and simply connected, it is sufficient to take $r$ above amongst the finite number of fundamental representations of ${}^{L}G$, and to twist by irreducible representations $\phi$ of ${\mathrm{WD}}_{K}$ of dimension at most $d$, where $d:=\lfloor N/2\rfloor$ for $N$ the dimension of any fundamental representation of highest dimension.
@@
For $G={\mathrm{SO}}_{4}(F)$, Theorem 6.3 is in agreement with the main result of [YZ] via the generic Langlands correspondence.
We now check that Conjecture 6.1 holds for stable discrete parameters of unitary groups. We will observe after the easy proof of this fact, that one can actually deduce from Theorem 6.3 (which is for split groups) a stronger result. Hence Conjecture 6.1 for non-split quasi-split groups, if correct, is in general not optimal, and maybe not very interesting. If $E/F$ is a separable quadratic extension, and $G:={\mathrm{U}}_{n}(E/F)$ is the corresponding quasi-split group, we call an admissible homomorphism from ${\mathrm{WD}}_{F}$ to ${}^{L}G$ stable discrete if its restriction to ${\mathrm{WD}}_{E}$ (interpreted as semi-simple $n$-dimensional representation) is irreducible. Such a parameter is automatically generic.
Proposition 6.6.
Let $n$ be a positive integer, $E/F$ be a separable quadratic extension, and $G:={\mathrm{U}}_{n}(E/F)$ the corresponding quasi-split group. Then Conjecture 6.1 holds for stable discrete homomorphisms from ${\mathrm{WD}}_{F}$ to ${}^{L}G$.
Proof.
[…]
$r_{0}:=\mathrm{Ind}_{{\mathrm{GL}}_{n}({\mathbb{C}})}^{{\mathrm{GL}}_{n}({\mathbb{C}})\rtimes{\mathbb{Z}}/2{\mathbb{Z}}}(\mathrm{St}).$
This irreducible representation of ${}^{L}G$ (obtained by inflating that of its quotient ${\mathrm{GL}}_{n}({\mathbb{C}})\rtimes{\mathbb{Z}}/2{\mathbb{Z}}$) is enough for our modest purpose. Under our assumption, the restriction of $(\phi_{i})_{E}$ to of $\phi_{i}$ to ${\mathrm{WD}}_{E}$ is irreducible. From this it follows that $(r_{0}\circ\phi_{i})_{E}=(\phi_{i})_{E}\oplus\alpha\circ(\phi_{i})_{E}$, hence either $(\phi_{2})_{E}=(\phi_{1})_{E}$, or $(\phi_{2})_{E}=\alpha\circ(\phi_{1})_{E}$. We deduce that $\phi_{2}=\phi_{1}$, or $\phi_{2}=\alpha\circ\phi_{1}$ thanks to for example to [Pra, Proposition 7.10]. ∎
Remark 6.7.
As previously mentioned, one can do better using Theorem 6.3. See Theorem 8.2 below, Case (a).
7. Unacceptable parameters
In this section we give an explicit example of a split group $G$, with unacceptable Langlands dual group, for which the statement of Conjecture 6.1 does not hold. First we introduce the notion of unacceptable homomorphism.
Definition 7.1.
Let $\widehat{G}$ be a complex connected reductive algebraic group, and $\varphi:\mathrm{W}_{F}\to\widehat{G}$ a semi-simple homomorphism. We say that $\varphi$ is acceptable if its local conjugacy class is equal to its global conjugacy class. Otherwise we say that $\varphi$ is unacceptable.
In particular, Theorem 5.3 says that if $\widehat{G}$ is acceptable, then any semi-simple homomorphism $\varphi:\mathrm{W}_{F}\to\widehat{G}$ is acceptable. Then we have the following proposition.
Proposition 7.2.
Suppose that $G$ is an $F$-split connected reductive group, and that there exists $\varphi_{1}:\mathrm{W}_{F}\to\widehat{G}$ which is unacceptable. Let $\varphi_{2}$ be locally conjugate but not globally conjugate to $\varphi_{1}$. Now take admissible homomorphisms $\phi_{1},\phi_{2}:\mathrm{W}_{F}\to{}^{L}G$ such that the semi-simple part of the Weil-Deligne parameter attached to $\phi_{1}$ is $\varphi_{1}$, and that of the Weil-Deligne parameter attached to $\phi_{2}$ is $\varphi_{2}$. Then for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
Proof.
For each $r$ the representations $r\circ\varphi_{1}$ and $r\circ\varphi_{2}$ are locally conjugate. Applying [Bou, Corollary of Proposition 6, a)] to the algebra of smooth compactly supported functions on $\mathrm{W}_{F}$, we deduce that $r\circ\varphi_{1}$ and $r\circ\varphi_{2}$ are globally conjugate, as their trace characters are equal. The equality of gamma factors follows. ∎
Now we deduce from [Yu, Example 2.1] due to Weidner ([Wei]), the following statement. It is interesting to compare it with the remark ending the introduction of [YZ].
Proposition 7.3.
Suppose that $14|q-$ 1. Then there exist non equivalent generic homomorphisms $\phi_{1},\phi_{2}:{\mathrm{WD}}_{F}\to{}^{L}{\mathrm{SO}}_{6}(F)$ such that for any irreducible representation $r$ of ${}^{L}G$, and any irreducible representation $\phi$ of ${\mathrm{WD}}_{F}$, one has
Proof.
Under the condition that $14|q-$ 1, and by class field theory, the group $(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}=\langle\gamma_{1},\ \gamma_{2}\rangle$ is a quotient of $\mathrm{W}_{F}$ because it is a quotient of $F^{\times}$. In particular let $\widetilde{\phi_{1}}:(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}\to{\mathrm{SL}}_{4}({\mathbb{C}})$ and $\widetilde{\phi_{2}}:(\frac{{\mathbb{Z}}}{4{\mathbb{Z}}})^{2}\to{\mathrm{SL}}_{4}({\mathbb{C}})$ the homomorphisms defined in [Yu, Example 2.1]: $\widetilde{\phi_{1}}(\gamma_{1})=\mathrm{diag}(1,1,i,-i)$, $\widetilde{\varphi_{1}}(\gamma_{2})=\mathrm{diag}(1,i,1,-i)$, and $\widetilde{\varphi_{2}}:=\overline{\widetilde{\varphi_{1}}}$. Set $\varphi_{i}$ to be the composition of $\widetilde{\phi_{i}}$ with the projection from ${\mathrm{SL}}_{4}({\mathbb{C}})$ to ${\mathrm{SL}}_{4}({\mathbb{C}})/\{\pm I_{4}\}\simeq{\mathrm{SO}}_{6}({\mathbb{C}})\subseteq{}^{L}{\mathrm{SO}}_{6}(F)$, and identify $\varphi_{i}$ with a semi-simple homomorphism $\mathrm{W}_{F}\to{}^{L}{\mathrm{SO}}_{6}(F)$ by inflation. One checks that the centralizer of both $\varphi_{1}$ and $\varphi_{2}$ are equal: seen in ${\mathrm{SL}}_{4}({\mathbb{C}})/\{\pm I_{4}\}$ they are equal to $D/\{\pm I_{4}\}$ where $D$ is the diagonal torus of ${\mathrm{SL}}_{4}({\mathbb{C}})$. In particular, according to [DHKM, Proposition 6.10], we can take a nilpotent element $N$ in the Lie algebra of ${\mathrm{SO}}_{6}({\mathbb{C}})$ such that $(\varphi_{1},N)$ and $(\varphi_{2},N)$ are Weil-Deligne parameters, which are moreover generic. We then take $\phi_{i}$ corresponding to $(\varphi_{i},N)$, and apply Proposition 7.2. ∎
An interesting question is to know if one can find such examples whenever $\widehat{G}$ is unacceptable, for some non Archimedean local field $F$, or even for all. This has to do with the classification of unacceptable homomorphisms.
@@ 8. A variant
Statement 8.1.
Let $G=\mathrm{G}(F)$ be a quasi-split group. Then there exists $K/F$ a finite extension and $r:{}^{L}\mathrm{G}(K)\to{\mathrm{GL}}(V)$ a faithful representation of dimension $N$, such that if $\phi_{i}:{\mathrm{WD}}_{F}\to{}^{L}G$ is a generic homomorphism for $i=1,2$, and
@@
Statement 8.1 is true for the following list of $G$, $K$ and $r$.
@@
    $G={\mathrm{GL}}_{n}(F)$, $K=F$ and $r:{\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{F}\to{\mathrm{GL}}_{n}({\mathbb{C}})$ is the identity on ${\mathrm{GL}}_{n}({\mathbb{C}})$ extended trivially on $\mathrm{W}_{F}$, or $G={\mathrm{U}}_{n}(E/F)$, $K=E$ a separable quadratic extension of $F$, and $r:{\mathrm{GL}}_{n}({\mathbb{C}})\times\mathrm{W}_{E}\to{\mathrm{GL}}_{n}({\mathbb{C}})$ is the identity on ${\mathrm{GL}}_{n}({\mathbb{C}})$ extended trivially on $\mathrm{W}_{E}$.
@@
The statement is obviously true in the first subcase of Case (a), thanks to Theorem 3.3 and Lemma 2.1. For unitary groups we recall that $\varphi_{1}\sim\varphi_{2}$ if and only if ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$ ([Pra, Proposition 7.10]). On the other hand, under the assumptions of Statement 8.1, we deduce from Theorem 3.3 that ${\varphi_{1}}_{|\mathrm{W}_{E}}\sim{\varphi_{2}}_{|\mathrm{W}_{E}}$, hence the result now follows from Lemma 2.1. In Case (b), the normalizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is equal to ×${\mathbb{C}}^{\times}\mathrm{G}_{2}({\mathbb{C}})$. This follows for example from applying [Gri, Theorem 1.1], and observing that the centralizer of $r(\mathrm{G}_{2}({\mathbb{C}}))$ is the center of ${\mathrm{GL}}_{7}({\mathbb{C}})$ according to Schur’s lemma. Then the theorem in Case (b) follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 1.1]. In cases (c), (d), (e) and (f), the normalizers are respectively ×${\mathbb{C}}^{\times}{\mathrm{SO}}_{2n+1}({\mathbb{C}})$, ×${\mathbb{C}}^{\times}{\mathrm{Sp}}_{2n}({\mathbb{C}})$ and ×${\mathbb{C}}^{\times}{\mathrm{O}}_{2n+2}({\mathbb{C}})$, as follows for example from [Gri, Section 2]. The result that we claim follows from Lemma 2.1, Theorem 3.3 and [Gri, Theorem 2.3]. ∎
@@
To conclude this paper, we make the following observation for classical groups when $F$ has characteristic zero. Assume the validity of the Gross-Prasad and Rallis conjecture and that the generic local Langlands correspondence $i$ of [JL3] coincides with that obtained by endoscopic transfer in [Art] and [Mok]. This latter fact might be known, and we refer to [AHKO, Theorem B.2 and Remark B.3] in the case of symplectic and split orthogonal groups. Then using [JL3, Theorem 6.20] and thanks to the uniqueness of generic representations inside L-packets with respect to a prescribed non degenerate character proved in [Ato], Theorem 8.2 coincides with the local converse theorems mentioned in the introduction, using Langlands-Shahidi gamma factors.